If G1,G2,.........G2n+1 are in GP,  the middle term is 50 the G1.G2n+1 is

Grade:

1 Answers

Arun
4 years ago
 
Dear student
 
Gn+1 = 50
 
a r^n = 50
 
now
 
G1* G 2n+1 = a * a* r^2n = a^2 * r*2n = 50^2 = 2500
 
hope it helps
 
Rwgards
Arun